OpenStudy (anonymous):

The length of a particular rectangle is 4 greater than the width of the rectangle. If the perimeter of the rectangle is 16, what is the area of the rectangle? 2 6 8 12 16

OpenStudy (sheraz12345):

Is this 4 times greater?

OpenStudy (imstuck):

Here's your drawing for this, using the perimeter formula:

OpenStudy (imstuck):



OpenStudy (imstuck):

If the length is 4 greater than the width, then it looks like this:

OpenStudy (imstuck):



OpenStudy (imstuck):

So fitting that into your formula, knowing that the perimeter measurement is 16, you get this:

OpenStudy (imstuck):

16 = 2W + 2(4+W) 16 = 2W + 8 + 2W 16 = 4W + 8 16 - 8 = 4W 8 = 4W 2 = W

OpenStudy (imstuck):

So the width is 2 and the length is 4 + 2 = 6. The area formula for a rectangle is L * W, so 2 * 6 = 12

OpenStudy (imstuck):

the fourth choice down is what you want.

OpenStudy (imstuck):

See that?